OpenStudy (safari321):

72 is what percent of 480?

OpenStudy (jeyekomon):

72/480*100 = 15 15 percent

hero (hero):

is/of = %/100 72/480 = x/100 480x = 7200 x = 7200/480 x = 15

OpenStudy (safari321):

thanks,that helps :)

hero (hero):

yup, anytime

OpenStudy (safari321):

:)